OpenStudy (anonymous):

given f(x) = x^2 and g(x) = square root of x, find g(f(x))

OpenStudy (anonymous):

\[g(f(x))=g(x^2)=\sqrt{x^2}\] How else can you write this?

OpenStudy (anonymous):

Close... What happens if \(x\) is negative?

OpenStudy (anonymous):

 Let's suppose the simplified form is indeed \(x\)...

OpenStudy (anonymous):

That means you would be able to plug in negative values of \(x\):